# Tri-resonant leptogenesis from modular symmetry neutrino models

### v1

@@ 1 Introduction
Remarkably, the seesaw model also provides an attractive explanation (which is known as the leptogenesis mechanism [10, 11]) for the baryon-antibaryon asymmetry of the Universe [12]
$Y_{\rm B}\equiv\frac{n_{\rm B}-n_{\rm\bar{B}}}{s}\simeq(8.69\pm 0.04)\times 10^{-11}\;,$ (7)
where $n_{\rm B}$ ($n_{\rm\bar{B}}$) denotes the baryon (antibaryon) number density and $s$ the entropy density. The leptogenesis mechanism works in a way as follows: a lepton-antilepton asymmetry $Y_{\rm L}\equiv(n_{\rm L}-n_{\rm\bar{L}})/s$ is firstly generated from the out-of-equilibrium and CP-violating decays of the right-handed neutrinos and then partly converted into the baryon-antibaryon asymmetry via the sphaleron processes: $Y_{\rm B}\simeq-cY_{\rm L}$ with $c=28/79$ or $8/23$ in the SM or MSSM (Minimal Supersymmetric Standard Model) framework. As is known, in the scenario that the right-handed neutrino masses are hierarchial there exists a lower bound about $10^{9}$ GeV for the right-handed neutrino mass scale in order to successfully reproduce the observed value of $Y_{\rm B}$ [13]. Obviously, such a high right-handed neutrino mass scale has no chance to be directly accessed by forseeable collider experiments. Fortunately, in the resonant leptogenesis scenario which is realized for nearly degenerate right-handed neutrinos [14], a successful leptogenesis can be realized even with TeV-scale right-handed neutrinos which have the potential to be directly accessed by running or upcoming experiments [15].
However, for some typical modular symmetry neutrino models (see the models provided in sections 3-5), the right-handed neutrino mass matrix takes a form as
$M_{\rm R}=\begin{pmatrix}M_{0}&0&0\\ 0&0&M_{0}\\ 0&M_{0}&0\end{pmatrix}\;.$ (8)
It can be diagonalized by the following unitary matrix
$U_{\rm R}=\begin{pmatrix}0&1&0\\ \frac{\rm i}{\sqrt{2}}&0&\frac{1}{\sqrt{2}}\\ -\frac{\rm i}{\sqrt{2}}&0&\frac{1}{\sqrt{2}}\end{pmatrix}\;,$ (9)
to give three degenerate right-handed neutrino masses (i.e., $M_{0}$). In this case, unfortunately, the leptogenesis mechanism is prohibited to work. But just as a coin has two sides, this kind of models can serve as a unique basis for realizations of the tri-resonant leptogenesis scenario: if the right-handed neutrino masses receive certain corrections so that their degeneracy is lifted to an appropriate extent, then the tri-resonant leptogenesis scenario will be naturally realized For a realization of resonant leptogenesis in the minimal seesaw model with a modular symmetry, see Ref. [16]. .
In this paper, we consider two minimal ways to generate the desired small mass splittings among the three right-handed neutrinos: one way is to modify $M_{\rm R}$ to a form as
$M_{\rm R}=\begin{pmatrix}M_{0}&0&0\\ 0&\mu&M_{0}\\ 0&M_{0}&\mu\end{pmatrix}\;.$ (10)
with $\mu\ll M_{0}$. Such a modification of $M_{\rm R}$ has the merit that it only introduces one additional parameter and it still can be diagonalized by the unitary matrix $U_{\rm R}$ in Eq. (9) (i.e., it does not alter the original neutrino mixing pattern). But the three right-handed neutrino masses are modified to $M_{0}-\mu$, $M_{0}$ and $M_{0}+\mu$, respectively. Another way is to consider the renormalization-group corrections for the right-handed neutrino masses [17]: if the energy scale $\Lambda_{\rm FS}$ of the flavor-symmetry physics that shapes the special textures of neutrino mass matrices is much higher than the right-handed neutrino mass scale $M_{0}$ where leptogenesis takes place (such a scenario is very likely given that the idea of modular flavor symmetry is just inspired from the string theory [7]), then the renormalization group evolution (RGE) effect between the scales of $\Lambda_{\rm FS}$ and $M_{0}$ may induce some mass splittings among the three right-handed neutrinos (see section 2.2). Such a correction of $M_{\rm R}$ has the following two merits: 1) this effect is spontaneous, provided that there is a considerable gap between the flavor-symmetry scale and the right-handed neutrino mass scale; 2) this effect is minimal, in the sense that it does not need to introduce additional parameters.
The remaining parts of this paper are organized as follows. In section 2 we will first give some basic formulas about resonant leptogenesis and RGE effects that will be used in our study. Then, we perform the study for three representative modular symmetry models which contain three degenerate right-handed neutrinos. We first consider the modular ${\rm A}_{4}$ symmetry model in section 3, then the modular ${\rm S}_{4}$ symmetry model in section 4, and finally the modular ${\rm A}_{5}$ symmetry model in section 5. In section 6, we summarize our main results.
2 Resonant leptogenesis and RGE effects
2.1 Some basics about resonant leptogenesis
As is known, if the temperature $T$ of the Universe at which leptogenesis takes place (approximately the right-handed neutrino mass scale) is below $10^{9}$ GeV, all the three lepton flavors have become distinguishable from one another and the lepton asymmetries stored in them should be tracked separately [18]. Such a three-flavor regime applies to the scenario studied in this paper (i.e., low-scale tri-resonant leptogenesis). In this case, the final lepton asymmetry can be approximately expressed as
$Y_{\rm L}\simeq r(\varepsilon_{e}\kappa_{e}+\varepsilon_{\mu}\kappa_{\mu}+\varepsilon_{\tau}\kappa_{\tau})\;.$ (11)
Here the factor $r\simeq 4\times 10^{-3}$ measures the ratio of the equilibrium number density of right-handed neutrinos to the entropy density. And $\varepsilon_{e}$ is a sum over three right-handed neutrinos (i.e., $\varepsilon_{e}=\varepsilon_{1e}+\varepsilon_{2e}+\varepsilon_{3e}$) of the following CP asymmetries:
$\varepsilon_{I\alpha}\equiv\frac{\Gamma\left(N_{I}\to L_{\alpha}+H\right)-\Gamma\left(N_{I}\to\overline{L}_{\alpha}+\overline{H}\right)}{\sum_{\alpha}\Gamma\left(N_{I}\to L_{\alpha}+H\right)+\Gamma\left(N_{I}\to\overline{L}_{\alpha}+\overline{H}\right)}\;,$ (12)
whose explicit expressions will be given below. Finally, $\kappa_{\alpha}$ are the efficiency factors (which are smaller than 1) due to the washout effects. Their concrete values can be calculated by numerically solving relevant Boltzmann equations, whose explicit expressions will also be given below.
To be explicit, in the basis where both the right-handed neutrino matrix and charged lepton mass matrix are diagonal, the CP asymmetries $\varepsilon_{I\alpha}$ defined in Eq. (12) can be expressed as [14]
$\varepsilon_{I\alpha}=\frac{\left|(\bar{Y}_{+})_{I\alpha}\right|^{2}-\left|(\bar{Y}_{-})_{I\alpha}\right|^{2}}{\left(\bar{Y}_{+}\bar{Y}^{\dagger}_{+}\right)_{II}+\left(\bar{Y}_{-}\bar{Y}^{\dagger}_{-}\right)_{II}}\;.$ (13)
Here $(\bar{Y}_{+})_{I\alpha}$ are the effective neutrino Yukawa couplings in the case of there existing three nearly degenerate right-handed neutrinos. They can be expressed as
$\left(\bar{Y}_{+}\right)_{I\alpha}=Y_{I\alpha}-{\rm i}V_{I\alpha}+{\rm i}\sum_{J,K=1}^{3}\left|\epsilon_{IJK}\right|Y_{J\alpha}$
×$\times\frac{M_{I}\left(\mathcal{M}_{IIJ}+\mathcal{M}_{JJI}\right)+{\rm i}R_{IK}\left[\mathcal{M}_{IKJ}\left(\mathcal{M}_{IIK}+\mathcal{M}_{KKI}\right)+\mathcal{M}_{JJK}\left(\mathcal{M}_{IKI}+\mathcal{M}_{KIK}\right)\right]}{M^{2}_{I}-M^{2}_{J}-2{\rm i}M^{2}_{I}A_{JJ}-2{\rm i}{\rm Im}(R_{IK})\left[M^{2}_{I}\left|A_{JK}\right|^{2}+M_{J}M_{K}{\rm Re}(A^{2}_{JK})\right]}\;,$ (14)
[…]
$R_{IJ}=\frac{M^{2}_{I}}{M^{2}_{I}-M^{2}_{J}-2{\rm i}M^{2}_{I}A_{JJ}}\;,$ (15)
with $f(x)=\sqrt{x}[1-(1+x)\ln(1+1/x)]$ being the loop function. And the corresponding CP-conjugate effective Yukawa couplings $\left(\bar{Y}_{-}\right)_{I\alpha}$ can be obtained by simply making the replacement $Y_{I\alpha}\to Y^{*}_{I\alpha}$ in the above expressions. Furthermore, in terms of these effective Yukawa couplings, the decay widths of right-handed neutrinos can be expressed as
$\Gamma\left(N_{I}\to L_{\alpha}+H\right)=\frac{M_{I}}{16\pi}\left|\left(\bar{Y}_{+}\right)_{I\alpha}\right|^{2}\;,\hskip 28.45274pt\Gamma\left(N_{I}\to\overline{L}_{\alpha}+\overline{H}\right)=\frac{M_{I}}{16\pi}\left|\left(\bar{Y}_{-}\right)_{I\alpha}\right|^{2}\;,$ (16)
and their branching ratios are given by
$B_{I\alpha}=\frac{\left|(\bar{Y}_{+})_{I\alpha}\right|^{2}+\left|(\bar{Y}_{-})_{I\alpha}\right|^{2}}{\left(\bar{Y}_{+}\bar{Y}^{\dagger}_{+}\right)_{II}+\left(\bar{Y}_{-}\bar{Y}^{\dagger}_{-}\right)_{II}}\;.$ (17)
In order to generate a nonzero lepton asymmetry in the Universe, one needs not only the above CP asymmetries, but also a departure from thermal equilibrium. In the latter processes, the evolutions of the number densities of the right-handed neutrinos $n_{N_{I}}$ and flavored lepton asymmetries $n_{L_{\alpha}}$ (which are normalized by the photon number density $n_{\gamma}$ to $\eta_{N_{I}}\equiv n_{N_{I}}/n_{\gamma}$ and $\eta_{L_{\alpha}}\equiv n_{L_{\alpha}}/n_{\gamma}$, respectively) can be tracked by the following Boltzmann equations [14]
$\frac{{\rm d}\eta_{N_{I}}}{{\rm d}z}$ $=$ $\frac{z}{H(z=1)}\Bigg{[}\left(1-\frac{\eta_{N_{I}}}{\eta_{N_{I}}^{\mathrm{eq}}}\right)\sum_{\alpha=e,\mu,\tau}\left(\Gamma^{D(I\alpha)}+\Gamma_{\text{Yukawa }}^{S(I\alpha)}+\Gamma_{\text{Gauge }}^{S(I\alpha)}\right)$
$-\frac{2}{3}\sum_{\alpha=e,\mu,\tau}\eta_{L_{\alpha}}\varepsilon_{I\alpha}\left(\hat{\Gamma}^{D(I\alpha)}+\hat{\Gamma}_{\text{Yukawa }}^{S(I\alpha)}+\hat{\Gamma}_{\text{Gauge }}^{S(I\alpha)}\right)\Bigg{]}\;,$
$\frac{{\rm d}\eta_{L_{\alpha}}}{{\rm d}z}$ $=$ $\frac{z}{H(z=1)}\Bigg{\{}\sum^{3}_{I=1}\varepsilon_{I\alpha}\left(\frac{\eta_{N_{I}}}{\eta_{N_{I}}^{\mathrm{eq}}}-1\right)\sum_{\beta=e,\mu,\tau}\left(\Gamma^{D(I\beta)}+\Gamma_{\text{Yukawa }}^{S(I\beta)}+\Gamma_{\text{Gauge }}^{S(I\beta)}\right)$ (18)
$-\frac{2}{3}\eta_{L_{\alpha}}\Bigg{[}\sum^{3}_{I=1}B_{I\alpha}\left(\tilde{\Gamma}^{D(I\alpha)}+\tilde{\Gamma}_{\text{Yukawa }}^{S(I\alpha)}+\tilde{\Gamma}_{\text{Gauge }}^{S(I\alpha)}+\Gamma_{\text{Yukawa }}^{W(I\alpha)}+\Gamma_{\text{Gauge }}^{W(I\alpha)}\right)$
$\hskip 42.67912pt+\sum_{\beta=e,\mu,\tau}\left(\Gamma_{\text{Yukawa }}^{\Delta L=2(\alpha\beta)}+\Gamma_{\text{Yukawa }}^{\Delta L=0(\alpha\beta)}\right)\Bigg{]}$
$-\frac{2}{3}\sum_{\beta=e,\mu,\tau}\eta_{L_{\beta}}\Bigg{[}\sum^{3}_{I=1}\varepsilon_{I\alpha}\varepsilon_{I\beta}\left(\Gamma_{\text{Yukawa }}^{W(I\beta)}+\Gamma_{\text{Gauge }}^{W(I\beta)}+\Gamma_{\text{Yukawa }}^{\Delta L=2(\alpha\beta)}-\Gamma_{\text{Yukawa }}^{\Delta L=0(\alpha\beta)}\right)\Bigg{]}\Bigg{\}}\;,$
where $z=M_{1}/T$ has been defined, and
$H\left(z\right)=\sqrt{\frac{4\pi^{3}g_{*}}{45}}\frac{M^{2}_{1}}{z^{2}M_{\mathrm{Pl}}}\;,$ (19)
is the Hubble parameter with $M_{\mathrm{Pl}}=1.221\times 10^{19}\mathrm{GeV}$ and $g_{*}\simeq 106.75$ or 228.75 (in the SM and MSSM frameworks, respectively) being the number of relativistic degrees of freedom in the early Universe. And the analytical expressions of the $\Gamma$ terms on the right-hand side of Eq. (18) can be found in Ref. [14]. Note that at the present epoch one has the relation $s\simeq 7.04n_{\gamma}$ for the entropy and photon number densities, and correspondingly $Y_{\rm L}\simeq\eta_{\rm L}/7.04$ (with $\eta_{\rm L}=\eta_{L_{e}}+\eta_{L_{\mu}}+\eta_{L_{\tau}}$).
2.2 Some basics about RGE effects
@@
As mentioned in the introduction section, the flavor-symmetry models which contain three degenerate right-handed neutrinos serve as a unique basis for realizations of the tri-resonant leptogenesis scenario after the right-handed neutrino masses receive certain corrections so that their degeneracy is lifted to an appropriate extent. And the renormalization group evolution effects from high to low energies can play the role of generating the desired right-handed neutrino mass splitings. In the literature, most of the flavor-symmetry models have been formulated in the MSSM framework for the following reason: in order to break the flavor symmetry in a proper way, one needs to introduce some flavon fields which transform as multiplets of the flavor symmetry and develop particular VEV alignments; and the most popular and perhaps natural approach to derive the desired flavon VEV alignments is provided by the so-called F-term alignment mechanism which is realized in a supersymmetric setup [6]. In particular, all of the modular flavor-symmetry models have been formulated in the MSSM framework since the modular flavor symmetry needs supersymmetry to preserve the holomorphicity of the modular form [7]. For these reasons, we will perform our study in the MSSM framework.
In the MSSM framework, the one-loop RGE evolution of the right-handed neutrino mass matrix and the Dirac neutrino Yukawa matrix $Y$ in the basis where $M_{\rm R}$ is diagonal are given by [19]
@@
where $t=\ln\left(M_{0}/\mu\right)$ has been defined, with $M_{0}$ and $\mu$ denoting the right-handed neutrino mass scale and renormalization scale, respectively. $\alpha=3g^{2}_{2}+\frac{3}{5}g^{2}_{1}-{\rm Tr}\left(3Y^{\dagger}_{\rm U}Y_{\rm U}+Y^{\dagger}Y\right)$, $g_{2}$ and $g_{1}$ are the $\rm SU\left(2\right)_{\rm L}$ and $\rm U\left(1\right)_{\rm Y}$ gauge coupling constants, and $Y_{{\rm U},l}$ are the Yukawa matrices for up-type quarks and charged leptons. $T$ is a $3\times 3$ anti-Hermitian matrix and its entries are given by
$T_{II}=0\;,$

### v3

@@ 2.1 Some formulas about leptogenesis
In order to generate a nonzero lepton asymmetry in the Universe, one needs not only the CP asymmetries, but also a departure from thermal equilibrium. In most cases, classical Boltzmann equations are sufficient for the calculation of the final asymmetry [11]. However, when lepton flavour effects are taken into account [16], different sets of classical Boltzmann equations apply depending whether the asymmetry is generated in the unflavored regime (where none of the charged-lepton Yukawa interactions has entered into equilibrium), in the two-flavour regime (where only the tauonic Yukawa interaction has entered into equilibrium), or in the three-flavour regime (where both the tauonic and muonic Yukawa interactions have entered into equilibrium). In particular, classical Boltzmann equations would fail in reproducing the correct result in the transition regimes between any two different flavor regimes. As our study aims to explore the full parameter space of the right-handed neutrino masses, we will employ the density matrix formalism [17] which applies to different flavor regimes as well as the transition regimes. In this formalism the evolutions of the right-handed neutrino number density $n_{N_{I}}$ and lepton asymmetry number density $n_{\alpha\beta}$ as functions of the temperature (via $z\equiv M_{1}/T$) can be tracked by the following equations [17]
$\frac{{\rm d}n_{N_{I}}}{{\rm d}z}$ $=$ $-D_{I}\left(n_{N_{I}}-n^{\rm eq}_{N_{I}}\right)\;,$
$\frac{{\rm d}n_{ee}}{{\rm d}z}$ $=$ $\sum_{I=1}^{3}\left\{\varepsilon^{\left(I\right)}_{ee}D_{I}\left(n_{N_{I}}-n^{\rm eq}_{N_{I}}\right)-W_{I}\left[\left|C^{0}_{Ie}\right|^{2}n_{ee}+{\rm Re}\left(C^{0}_{Ie}C^{0}_{I\mu}n^{*}_{\mu e}+C^{0}_{Ie}C^{0}_{I\tau}n^{*}_{\tau e}\right)\right]\right\}\;,$
$\frac{{\rm d}n_{\mu\mu}}{{\rm d}z}$ $=$ $\sum_{I=1}^{3}\left\{\varepsilon^{\left(I\right)}_{\mu\mu}D_{I}\left(n_{N_{I}}-n^{\rm eq}_{N_{I}}\right)-W_{I}\left[\left|C^{0}_{I\mu}\right|^{2}n_{\mu\mu}+{\rm Re}\left(C^{0}_{I\mu}C^{0}_{Ie}n_{\mu e}+C^{0}_{I\mu}C^{0}_{I\tau}n^{*}_{\tau\mu}\right)\right]\right\}\;,$
$\frac{{\rm d}n_{\tau\tau}}{{\rm d}z}$ $=$ $\sum_{I=1}^{3}\left\{\varepsilon^{\left(I\right)}_{\tau\tau}D_{I}\left(n_{N_{I}}-n^{\rm eq}_{N_{I}}\right)-W_{I}\left[\left|C^{0}_{I\tau}\right|^{2}n_{\tau\tau}+{\rm Re}\left(C^{0}_{I\tau}C^{0}_{Ie}n_{\tau e}+C^{0}_{I\tau}C^{0}_{I\mu}n_{\tau\mu}\right)\right]\right\}\;,$
$\frac{{\rm d}n_{\tau\mu}}{{\rm d}z}$ $=$ $\sum_{I=1}^{3}\bigg\{\varepsilon^{\left(I\right)}_{\tau\mu}D_{I}\left(n_{N_{I}}-n^{\rm eq}_{N_{I}}\right)-\frac{1}{2}W_{I}\bigg[n_{\tau\mu}\left(\left|C^{0}_{I\tau}\right|^{2}+\left|C^{0}_{I\mu}\right|^{2}\right)+C^{0*}_{I\mu}C^{0}_{I\tau}\left(n_{\tau\tau}+n_{\mu\mu}\right)$
$\hskip 28.45274pt+C^{0*}_{Ie}C^{0}_{I\tau}n^{*}_{\mu e}+C^{0*}_{I\mu}C^{0}_{Ie}n_{\tau e}\bigg]\bigg\}-\left[\frac{{\rm Im}\left(\Lambda_{\tau}\right)}{Hz}+\frac{{\rm Im}\left(\Lambda_{\mu}\right)}{Hz}\right]n_{\tau\mu}\;,$
$\frac{{\rm d}n_{\tau e}}{{\rm d}z}$ $=$ $\sum_{I=1}^{3}\bigg\{\varepsilon^{\left(I\right)}_{\tau e}D_{I}\left(n_{N_{I}}-n^{\rm eq}_{N_{I}}\right)-\frac{1}{2}W_{I}\bigg[n_{\tau e}\left(\left|C^{0}_{Ie}\right|^{2}+\left|C^{0}_{I\tau}\right|^{2}\right)+C^{0*}_{Ie}C^{0}_{I\tau}\left(n_{ee}+n_{\tau\tau}\right)$
$\hskip 28.45274pt+C^{0*}_{I\mu}C^{0}_{I\tau}n_{\mu e}+C^{0*}_{Ie}C^{0}_{I\mu}n_{\tau\mu}\bigg]\bigg\}-\frac{{\rm Im}\left(\Lambda_{\tau}\right)}{Hz}n_{\tau e}\;,$
$\frac{{\rm d}n_{\mu e}}{{\rm d}z}$ $=$ $\sum_{I=1}^{3}\bigg\{\varepsilon^{\left(I\right)}_{\mu e}D_{I}\left(n_{N_{I}}-n^{\rm eq}_{N_{I}}\right)-\frac{1}{2}W_{I}\bigg[n_{\mu e}\left(\left|C^{0}_{Ie}\right|^{2}+\left|C^{0}_{I\mu}\right|^{2}\right)+C^{0*}_{Ie}C^{0}_{I\mu}\left(n_{ee}+n_{\mu\mu}\right)$ (23)
$\hskip 28.45274pt+C^{0*}_{Ie}C^{0}_{I\tau}n^{*}_{\tau\mu}+C^{0*}_{I\tau}C^{0}_{I\mu}n_{\tau e}\bigg]\bigg\}-\frac{{\rm Im}\left(\Lambda_{\mu}\right)}{Hz}n_{\mu e}\;,$
where $C^{0}_{I\alpha}$ have been defined in Eq. (22). In these equations, $n^{\rm eq}_{N_{I}}$ are the equilibrium number density of three right-handed neutrinos. The decay (washout) parameters $D_{I}\left(W_{I}\right)$ are given by
$D_{I}\left(z\right)=K_{I}x_{I}z\frac{\mathcal{K}_{1}\left(z_{I}\right)}{\mathcal{K}_{2}\left(z_{I}\right)}\;,\qquad W_{I}\left(z\right)=\frac{1}{4}K_{I}\sqrt{x_{I}}\;\mathcal{K}_{1}\left(z_{I}\right)z^{3}_{I}\;,$ (24)
with $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$ being the modified Bessel functions of the second kind and
$x_{I}\equiv\frac{M^{2}_{I}}{M^{2}_{1}}\;,\qquad z_{I}\equiv z\sqrt{x_{I}}\;,\qquad K_{I}\equiv\frac{\Gamma_{I}}{H\left(T=M_{1}\right)}\;.$ (25)
Moreover, the Hubble expansion rate $H$ is given by
$H=\sqrt{\frac{4\pi^{3}g_{*}}{45}}\frac{M^{2}_{1}}{z^{2}M_{\rm Pl}}\;,$ (26)
in which $M_{\rm Pl}=1.22\times 10^{19}$GeV is the Planck mass and $g_{*}\simeq 106.75$ or 228.75 (in the SM and MSSM frameworks, respectively) is the number of relativistic degrees of freedom in the early Universe. Last but not least, $\Lambda_{\alpha}$ is the self-energy of $\alpha$-flavored leptons, and its imaginary part is given by
${\rm Im}\left(\Lambda_{\alpha}\right)\simeq 8\times 10^{-3}y^{2}_{\alpha}T\;,$ (27)
with $y_{\alpha}$ being the charged-lepton Yukawa coupling.
2.2 Some formulas about RGE effects
@@
As mentioned in the introduction section, the flavor-symmetry models which contain three degenerate right-handed neutrinos serve as a unique basis for realizations of the tri-resonant leptogenesis scenario after the right-handed neutrino masses receive certain corrections so that their degeneracy is lifted to an appropriate extent. And the RGE effects between the flavor-symmetry and leptogenesis scales can play the role of generating the desired right-handed neutrino mass splitings. In the literature, most of the flavor-symmetry models have been formulated in the MSSM framework for the following reason: in order to break the flavor symmetry in a proper way, one needs to introduce some flavon fields which transform as multiplets of the flavor symmetry and develop particular VEV alignments; and the most popular and perhaps natural approach to derive the desired flavon VEV alignments is provided by the so-called F-term alignment mechanism which is realized in a supersymmetric setup [5]. In particular, the modular flavor-symmetry models have been formulated in the MSSM framework since the modular flavor symmetry needs supersymmetry to preserve the holomorphicity of the modular form [7]. For these reasons, we will perform our study in the MSSM framework.
In the MSSM framework, the relevant renormalization-group equations have been given in Ref. [18]. First of all, the lepton fields can be transformed by a unitary transformation to the basis in which the right-handed neutrino mass matrix is diagonal. In this basis, the RGE evolution of three right-handed neutrino masses $M_{I}$ and the neutrino Yukawa coupling matrix $Y$ can be obtained by using the methods outlined in [14, 19]
@@
where $t=\ln\left(M_{0}/\mu\right)$ has been defined, with $M_{0}$ and $\mu$ denoting the right-handed neutrino mass scale and renormalization scale, respectively. And one has $\alpha=3g^{2}_{2}+\frac{3}{5}g^{2}_{1}-{\rm Tr}\left(3Y^{\dagger}_{u}Y_{u}+Y^{\dagger}Y\right)$, where $g_{2}$ and $g_{1}$ are the $\rm SU\left(2\right)_{\rm L}$ and $\rm U\left(1\right)_{\rm Y}$ gauge coupling constants, and $Y_{u,l}$ are the Yukawa coupling matrices for the up-type quarks and charged leptons. $T$ is a $3\times 3$ anti-Hermitian matrix and its non-zero entries are given by